# GUIDE: Enhancing Gradient Inversion Attacks in Federated Learning with Denoising Models

### v1

@@ 5.3 Attack Scenario 2: Face Recognition
In this scenario, we consider ROG Yue et al. (2023) as the base attack $\mathcal{A}$, using ResNet-18 He et al. (2016) as shared model, trained for a face recognition task. The victim client’s data consists of $128\times 128$ images from a subset of identities in VGGFace2 Cao et al. (2018), while the surrogate dataset consists of distinct and non-overlapping identities from the same dataset. The denoising techniques employed follow the methods described in Luo et al. (2023b, a).
@@
GUIDE Setup. The setup phase is performed by the attacker by sampling data from the surrogate dataset and computing model updates $\nabla^{\prime}$ using the FedAVG algorithm with 5 local epochs and a batch size of 16 and the learning rate $\eta=5\times 10^{-3}$, following the same setup of the original work Yue et al. (2023).
@@
As described in Section 2, ROG Yue et al. (2023) follows a two-step optimization process, first performing the attack in a reduced-dimensional space to simplify optimization and improve convergence, then using restoration modules to reconstruct images in the original high-dimensional space, enhancing perceptual quality.
@@
In this context, GUIDE enables the training of specialized restoration models tailored to the specific distortions introduced during the optimization in the reduced space, rather than relying on general-purpose restoration models as presented in the original paper Yue et al. (2023). To achieve this, the attacker collects images after the first step of the attack and constructs a denoising dataset, $D_{den}$. In our experiments, we gathered 106,796 images.
@@
For the denoising process, we employed a diffusion-based denoising model as $\mathcal{M}_{den}$. After collecting the data, we performed an 80%/20% split for training and testing and trained the model using the original hyperparameters from Luo et al. (2023b). The diffusion model, based on a Conditional NAFNet architecture Luo et al. (2023b), was trained for 700,000 iterations with a batch size of 4, using the Lion optimizer Chen et al. (2024) and optimizing the $L_{1}$ loss function. In this case, the model’s input consists of images from the reduced search space ($32\times 32$ pixels), while the output is in the original image space ($128\times 128$ pixels). Consequently, the denoising model is trained for a super-resolution task with a scaling factor of 4.
@@
Figure 5: Example of reconstructions obtained by different methods on 8 images in Scenario  2. The model update is computed over 16 images with a resolution of $128\times 128$ with ResNet-18 model. All the images are not included in any noisy image seen during the training of the denoising models. (a) Images reconstructed with ROG Yue et al. (2023). (b) Images reconstructed with ROG Yue et al. (2023) enhanced with GUIDE (ours). (c) Original images used to compute the gradient.
@@
Table 5: Performance comparison of attacks in Scenario 2 when client data are taken from VGGFace2 Cao et al. (2018). Updates are computed across 30 batches of 16 images, and results are averaged.
GUIDE Reconstruction. In this section, we assess the capacity of GUIDE to enhance image reconstruction of the ROG Yue et al. (2023) attack. Specifically, we evaluate the GUIDE, where the denoising model $M_{den}$ is applied at the end of the first optimization step of the attack, with $\mathcal{D}_{iters}=\{100\}$. This choice is motivated by the fact that $M_{den}$ is specifically trained to reconstruct images from the reduced space back to the original input space and, therefore, cannot be directly integrated into the optimization process conducted within the reduced space. Performances are compared with the original ROG configuration presented in Yue et al. (2023).
Table 5 presents a comparative analysis of reconstruction performances between ROG and GUIDE. The results show that GUIDE consistently outperforms ROG across all evaluated metrics. In particular, GUIDE achieves higher PSNR and SSIM values, indicating improved pixel-wise accuracy and structural similarity with the original images. Moreover, GUIDE achieves a lower LPIPS score, indicating superior perceptual quality, and significantly reduces the DreamSim score by 46%, demonstrating a stronger alignment with human perception. These quantitative improvements align with the visual results presented in Figure 5, where images reconstructed with GUIDE retain significantly more details than those produced by ROG. In some cases, the enhanced quality of GUIDE’s reconstructions may even enable face recognizability, raising concerns about potential privacy risks for the attacked clients.
@@
Table 6: Performance comparison of attacks in Scenario 2 when client data are taken from FFHQ Karras (2019). Updates are computed across 30 batches of 16 images, and results are averaged.
Robustness to Distribution Shift. A natural question that arises is whether the reliance on surrogate data limits the applicability of GUIDE. In this section, we evaluate the effectiveness of our method under distribution shift by employing a surrogate dataset a surrogate dataset that is not only distinct from the dataset used by the victim client, but is also sampled from an entirely different database. Specifically, the denoising model remains trained on the surrogate dataset described in Table 2 sampled from VGGFace2 Cao et al. (2018), but the victim client’s updates are computed on data from FFHQ Karras (2019), further increasing the distributional shift between the surrogate and client data. This serves as a robustness test to analyze how GUIDE performs when the attacker has access to a visually similar but distinct dataset for training the denoising model while the actual gradients originate from a separate data distribution. The results in Table 6 show that GUIDE continues to outperform ROG across all metrics, demonstrating its ability to enhance reconstruction quality even when attacker has access data that comes from a separate distribution with respect to those of victim client. In particular, the 32% reduction in DreamSim suggests that the reconstructed images retain a high degree of perceptual similarity to the original ones, indicating that privacy leakage remains significant despite the increased distributional shift between the client and surrogate datasets.
5.4 Defensive Measures
@@
Table 7: Performance comparison of GUIDE and CPA Kariyappa et al. (2023) attacks in Scenario 1 under differentially private noise perturbation. Results are averaged over 30 batches of 8 images for each metric and $\sigma$ value.
As GIAs advance, various defensive measures have been developed to protect client privacy. Since the client updates $\nabla$ shared during training is the primary target of GIAs, multiple strategies have been proposed to prevent attackers from accessing and exploiting this information. The defensive measures can be broadly categorized into cryptographic approaches, formal perturbation methods, and heuristic-based perturbations. Cryptographic techniques, such as Homomorphic Encryption (HE) Acar et al. (2018) and Secure Aggregation (SA) Bonawitz et al. (2017); Mohamad et al. (2023), can be integrated into FL training to ensure strong privacy guarantees by preventing direct access to updates. However, these mechanisms introduce significant computational and communication overhead Mohamad et al. (2023); Xie et al. (2024), limiting their practicality in resource-constrained FL deployments. An alternative strategy involves perturbing updates before they are shared. This can be achieved through Differential Privacy (DP) Ponomareva et al. (2023), which adds carefully calibrated noise to updates based on a rigorous mathematical framework, or with heuristic-based perturbation methods, which modify gradients in an ad-hoc manner to reduce the privacy risk. While these perturbation techniques may counter GIAs, they inherently impact model accuracy and utility due to the noise introduced into the training process.
In this section, we analyze the effectiveness of defensive measures and assess whether GUIDE remains effective in enhancing reconstruction quality when they are applied. Specifically, we apply privacy-preserving techniques to both attack scenarios considered in this work. To ensure a fair and consistent assessment, we carefully tune these defenses using the parameters established in CPA Kariyappa et al. (2023) and ROG Yue et al. (2023).
@@
Table 8: Performance comparison of ROG Yue et al. (2023) and GUIDE attacks in Scenario 2 under different gradient perturbation measures. Results are averaged over 30 batches.
@@
Attack Scenario 1: Image Classification with Differential Privacy. This set of experiments demonstrates that when the base attack $\mathcal{A}$ fails to reconstruct a meaningful noisy version of the images, GUIDE does not provide any improvement in reconstruction quality. Table 7 shows that GUIDE fails to enhance reconstruction quality under increasing differentially private noise levels. This is due to the limitations of the CPA attack itself: when CPA produces highly degraded reconstructions, the denoising model in GUIDE lacks sufficient structural information to restore meaningful details. As the noise $\sigma$ increases, both methods exhibit lower PSNR and SSIM, while LPIPS and DreamSim scores worsen, confirming that GUIDE cannot improve reconstruction quality when the base attack fails to retain recognizable image features. The $\sigma$ values used in our evaluation were chosen based on the parameters reported in Kariyappa et al. (2023), ensuring a fair and consistent comparison with prior work.
@@
Attack Scenario 2: Face Recognition with Gradient Perturbation Techniques. In this section, we demonstrate that when the base attack $\mathcal{A}$ produces meaningful noisy images, the GUIDE approach can effectively enhance reconstruction quality even in the presence of defensive measures, without altering the underlying threat model. By using ROG Yue et al. (2023) as the base attack, we assess the impact of the following gradient compression techniques, using the same parameter configuration of the original paper Yue et al. (2023): (i) QSGD Alistarh et al. (2017) with a 3-bit quantizer; and (ii) Top-k Aji and Heafield (2017) with a sparsity parameter of 0.95. To adapt to these defensive measures, the attacker follows the same procedure described in Section 5.3 to train two defense-specific denoising models using the surrogate dataset. The training pipeline remains unchanged, except that the dataset $D_{den}$ is now constructed by considering the effects of these defenses, ensuring that the noisy images reflect the perturbations applied on the client side. In our experiments, we collect 81,052 images by applying the QSGD defense and 76,485 images by applying the Top-k defense. To obtain specialized denoising models, $\mathcal{M}_{den}^{\text{qsgd}}$ and $\mathcal{M}_{den}^{\text{Top-{k}}}$, we fine-tune the model previously trained in Section 5.3, ensuring it is optimized for reconstructing images under the specific distortions introduced by each defensive measure.
@@
The results in Table 8 demonstrate that GUIDE consistently enhances reconstruction quality across both gradient perturbation defenses. For the Top-k defense, GUIDE achieves a 49.15% reduction in DreamSim, indicating that the reconstructed images are not only sharper but also more semantically aligned with human perception. This suggests an increased risk of privacy leakage despite the applied defense. For QSGD, GUIDE improves reconstruction quality across all evaluated metrics, though to a lesser extent than with Top-k. The 29.33% reduction in DreamSim suggests that even under aggressive quantization, GUIDE enhances perceptual similarity, making reconstructed images more recognizable. Overall, these findings indicate that while gradient compression techniques reduce reconstruction quality, they do not fully prevent privacy leakage. The ability of GUIDE to adapt by training defense-specific denoising models demonstrates that attackers can partially bypass these defenses, particularly when the base attack retains structural information about the original images.
6 Conclusion

### v2

@@ V-C Attack Scenario 2: Face Recognition
In this scenario, we consider ROG [15] as the base attack $\mathcal{A}$, using ResNet-18 [41] as the shared model, trained for a face recognition task. The victim client’s data consists of $128\times 128$ images from a subset of identities in VGGFace2 [34], while the surrogate dataset consists of distinct and non-overlapping identities from the same dataset. The denoising techniques employed follow the methods described in [32, 33].
@@
GUIDE Setup. The setup phase is performed by the attacker by sampling data from the surrogate dataset and computing model updates $\nabla^{\prime}$ using the FedAVG algorithm with 5 local epochs and a batch size of 16 and the learning rate $\eta=5\times 10^{-3}$, following the same setup of the original work [15].
@@
As described in Section II, ROG [15] follows a two-step optimization process, first performing the attack in a reduced-dimensional space to simplify optimization and improve convergence, then using restoration modules to reconstruct images in the original high-dimensional space, enhancing perceptual quality.
@@
In this context, GUIDE enables the training of specialized restoration models tailored to the specific distortions introduced during the optimization in the reduced space, rather than relying on general-purpose restoration models as presented in the original paper [15]. To achieve this, the attacker collects images after the first step of the attack and constructs a denoising dataset, $D_{den}$. In our experiments, we gathered 106,796 images.
@@
For the denoising process, we employed a diffusion-based denoising model as $\mathcal{M}_{den}$. After collecting the data, we performed an 80%/20% split for training and testing and trained the model using the original hyperparameters from [32]. The diffusion model, based on a Conditional NAFNet architecture [32], was trained for 700,000 iterations with a batch size of 4, using the Lion optimizer [43] and optimizing the $L_{1}$ loss function. In this case, the model’s input consists of images from the reduced search space ($32\times 32$ pixels), while the output is in the original image space ($128\times 128$ pixels). Consequently, the denoising model is trained for a super-resolution task with a scaling factor of 4.
@@
Figure 5: Example of reconstructions obtained by different methods on 8 images in Scenario  2. The model update is computed over 16 images with a resolution of $128\times 128$ with the ResNet-18 model. All the images are not included in any noisy image seen during the training of the denoising models. (a) Images reconstructed with ROG [15]. (b) Images reconstructed with ROG [15] enhanced with GUIDE (ours). (c) Original images used to compute the gradient.
GUIDE Reconstruction. In this section, we assess the capacity of GUIDE to enhance image reconstruction of the ROG [15] attack. Specifically, we evaluate the GUIDE, where the denoising model $M_{den}$ is applied at the end of the first optimization step of the attack, with $\mathcal{D}_{iters}=\{100\}$. This choice is motivated by the fact that $M_{den}$ is specifically trained to reconstruct images from the reduced space back to the original input space and, therefore, cannot be directly integrated into the optimization process conducted within the reduced space. Performances are compared with the original ROG configuration presented in [15].
Table V presents a comparative analysis of reconstruction performances between ROG and GUIDE. The results show that GUIDE consistently outperforms ROG across all evaluated metrics. In particular, GUIDE achieves higher PSNR and SSIM values, indicating improved pixel-wise accuracy and structural similarity with the original images. Moreover, GUIDE achieves a lower LPIPS score, indicating superior perceptual quality, and significantly reduces the DreamSim score by 46%, demonstrating a stronger alignment with human perception. These quantitative improvements align with the visual results presented in Figure 5, where images reconstructed with GUIDE retain significantly more details than those produced by ROG. In some cases, the enhanced quality of GUIDE’s reconstructions may even enable face recognizability, raising concerns about potential privacy risks for the attacked clients.
@@
TABLE V: Performance comparison of attacks in Scenario 2 when client data are taken from VGGFace2 [34]. Updates are computed across 30 batches of 16 images, and results are averaged.
Robustness to Distribution Shift. A natural question that arises is whether the reliance on surrogate data limits the applicability of GUIDE. In this section, we evaluate the effectiveness of our method under distribution shift by employing a surrogate dataset that is not only distinct from the dataset used by the victim client but is also sampled from an entirely different database. Specifically, the denoising model remains trained on the surrogate dataset described in Table II sampled from VGGFace2 [34], but the victim client’s updates are computed on data from FFHQ [45], further increasing the distributional shift between the surrogate and client data. This serves as a robustness test to analyze how GUIDE performs when the attacker has access to a visually similar but distinct dataset for training the denoising model while the actual gradients originate from a separate data distribution. The results in Table VI show that GUIDE continues to outperform ROG across all metrics, demonstrating its ability to enhance reconstruction quality even when the attacker has access to data that comes from a separate distribution with respect to those of the victim client. In particular, the 32% reduction in DreamSim suggests that the reconstructed images retain a high degree of perceptual similarity to the original ones, indicating that privacy leakage remains significant despite the increased distributional shift between the client and surrogate datasets.
@@
TABLE VI: Performance comparison of attacks in Scenario 2 when client data are taken from FFHQ [45]. Updates are computed across 30 batches of 16 images, and results are averaged.
V-D Defensive Measures
@@
TABLE VII: Performance comparison of GUIDE and CPA [20] attacks in Scenario 1 under differentially private noise perturbation. Results are averaged over 30 batches of 8 images for each metric and $\sigma$ value.
As GIAs advances, various defensive measures have been developed to protect client privacy. Since the client updates $\nabla$ shared during training are the primary target of GIAs, multiple strategies have been proposed to prevent attackers from accessing and exploiting this information. The defensive measures can be broadly categorized into cryptographic approaches, formal perturbation methods, and heuristic-based perturbations. Cryptographic techniques, such as Homomorphic Encryption (HE) [46] and Secure Aggregation (SA) [47, 48], can be integrated into FL training to ensure strong privacy guarantees by preventing direct access to updates. However, these mechanisms introduce significant computational and communication overhead [48, 49], limiting their practicality in resource-constrained FL deployments. An alternative strategy involves perturbing updates before they are shared. This can be achieved through Differential Privacy (DP) [37], which adds carefully calibrated noise to updates based on a rigorous mathematical framework, or with heuristic-based perturbation methods, which modify gradients in an ad-hoc manner to reduce the privacy risk. While these perturbation techniques may counter GIAs, they inherently impact model accuracy and utility due to the noise introduced into the training process.
In this section, we analyze the effectiveness of defensive measures and assess whether GUIDE remains effective in enhancing reconstruction quality when they are applied. Specifically, we apply privacy-preserving techniques to both attack scenarios considered in this work. To ensure a fair and consistent assessment, we carefully tune these defenses using the parameters established in CPA [20] and ROG [15].
@@
TABLE VIII: Performance comparison of ROG [15] and GUIDE attacks in Scenario 2 under different gradient perturbation measures. Results are averaged over 30 batches.
@@
Attack Scenario 1: Image Classification with Differential Privacy. This set of experiments demonstrates that when the base attack $\mathcal{A}$ fails to reconstruct a meaningful noisy version of the images, GUIDE does not provide any improvement in reconstruction quality. Table VII shows that GUIDE fails to enhance reconstruction quality under increasing differentially private noise levels. This is due to the limitations of the CPA attack itself: when CPA produces highly degraded reconstructions, the denoising model in GUIDE lacks sufficient structural information to restore meaningful details. As the noise $\sigma$ increases, both methods exhibit lower PSNR and SSIM, while LPIPS and DreamSim scores worsen, confirming that GUIDE cannot improve reconstruction quality when the base attack fails to retain recognizable image features. The $\sigma$ values used in our evaluation were chosen based on the parameters reported in [20], ensuring a fair and consistent comparison with prior work.
@@
Attack Scenario 2: Face Recognition with Gradient Perturbation Techniques. In this section, we demonstrate that when the base attack $\mathcal{A}$ produces meaningful noisy images, the GUIDE approach can effectively enhance reconstruction quality even in the presence of defensive measures, without altering the underlying threat model. By using ROG [15] as the base attack, we assess the impact of the following gradient compression techniques, using the same parameter configuration of the original paper [15]: (i) QSGD [38] with a 3-bit quantizer; and (ii) Top-k [39] with a sparsity parameter of 0.95. To adapt to these defensive measures, the attacker follows the same procedure described in Section V-C to train two defense-specific denoising models using the surrogate dataset. The training pipeline remains unchanged, except that the dataset $D_{den}$ is now constructed by considering the effects of these defenses, ensuring that the noisy images reflect the perturbations applied on the client side. In our experiments, we collect 81,052 images by applying the QSGD defense and 76,485 images by applying the Top-k defense. To obtain specialized denoising models, $\mathcal{M}_{den}^{\text{qsgd}}$ and $\mathcal{M}_{den}^{\text{Top-{k}}}$, we fine-tune the model previously trained in Section V-C, ensuring it is optimized for reconstructing images under the specific distortions introduced by each defensive measure.
@@
The results in Table VIII demonstrate that GUIDE consistently enhances reconstruction quality across both gradient perturbation defenses. For the Top-k defense, GUIDE achieves a 49.15% reduction in DreamSim, indicating that the reconstructed images are not only sharper but also more semantically aligned with human perception. This suggests an increased risk of privacy leakage despite the applied defense. For QSGD, GUIDE improves reconstruction quality across all evaluated metrics, though to a lesser extent than with Top-k. The 29.33% reduction in DreamSim suggests that even under aggressive quantization, GUIDE enhances perceptual similarity, making reconstructed images more recognizable. Overall, these findings indicate that while gradient compression techniques reduce reconstruction quality, they do not fully prevent privacy leakage. The ability of GUIDE to adapt by training defense-specific denoising models demonstrates that attackers can partially bypass these defenses, particularly when the base attack retains structural information about the original images.
VI Conclusion